OpenStudy (twalt13):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. 4-6+5-7+6-8 + ... + 4n( 4n + 2) = 4(4n+1)(8n+7)/6

OpenStudy (anonymous):

Could you rewrite that expression? I'm confused on what the inductive formula is.

OpenStudy (campbell_st):

can you rewrite the sum of the terms pls as it appears that it fails the test for n = 1 \[4(4 + 2) = \frac{4(4 + 1)(8 + 7)}{6}\]

OpenStudy (twalt13):

4-6+5-7+6-8

OpenStudy (anonymous):

I meant the expression to the farmost right of the equal sign. Is it what campbell_st posted?

OpenStudy (twalt13):

yes

OpenStudy (campbell_st):

what is the sum term... is it correct...? \[= \frac{4(4n + 1)(8n+7)}{6}\]

OpenStudy (twalt13):

yea there it is he got it

OpenStudy (campbell_st):

ok... so the 1st thing to so is prove it correct for n = 1 so the 1st term is equal to the sum of 1 term \[T_{1} = 4\times 1(4 \times 1 + 2) = 24\] now look at the sum when n = 1 \[S_{1} = \frac{4(4 \times 1 + 1)\times (8 \times 1 + 7)}{6} = 50\] so there would appear to be a mistake somewhere... if you are to prove it by induction

OpenStudy (twalt13):

so it is false

OpenStudy (campbell_st):

well it would appear to fail the test where n = 1

OpenStudy (twalt13):

Thanks

OpenStudy (twalt13):

wait what about k ? @campbell_st

OpenStudy (campbell_st):

well you assume its true for n = k so \[S_{k} = \frac{4(4k + 1)(8k + 7)}{6}\] but there is no use looking at n = k since it failed for n = 1